in this system of equations, find x: 2x+3y=6 & 3x+3y=9
Answer by mananth(16946)

2x+3y=6
3x+3y=9
..
subtract the equations
2x+3y-3x-3y=6-9
-x=-3
x=3
..
plug value of x in equation 1
6+3y=6
3y=0
y=0